InkSans352:

Pedro works as a newspaper deliverer. He receives a base pay plus an additional amount per newspaper delivered. Last week, Pedro delivered 5 newspapers and earned $37.50. This week, he delivered 20 newspapers and earned $75. Let x equal the number of newspapers Pedro delivers and y be the earnings he receives in dollars. Which linear function models the scenario?

InkSans352:

Options: g(x) = 0.4x + 35.5 g(x) = 0.4x + 67 g(x) = 2.5x + 25 g(x) = 2.5x + 50

Narad:

\[y=a+bx\]

InkSans352:

Thanks so much!